# Unexpected consequences of Post-Quantum theories in the graph-theoretical approach to correlations

### v1

@@ I Introduction
Quantum Theory (QT) is one of the most successful scientific theories ever created by humankind; in that its predictions match with great accuracy empirical data. One paramount example is the loophole-free confirmation of the existence of quantum Bell-nonlocality [1, 2, 3, 4, 5]. Nonetheless, it remains a challenge to formulate it as a consequence of physically justifiable principles (of Nature), similar to what occurs with Einstein’s Special Relativity. Several distinct approaches have been proposed to address this problem. For example, certain approaches adopt an axiomatic framework, beginning with a more physically intuitive set of axioms from which one aims to derive the standard and abstract postulates of quantum theory (the Hilbert space structure, density matrices and measurement operators acting on it, and Born’s rule) [6, 7, 8, 9]. Another program is the one started by Popescu and Rohrlich’s seminal work in trying to answer the question: are quantum correlations singled out by the non-signaling principle [10]? The answer was found to be negative, as there exist correlations that satisfy the non-signaling condition but cannot be reproduced within the framework of quantum theory, with Popescu-Rohrlich boxes being the most prominent example [11]. Although this approach does not fully explain quantum correlations, it raises the broader question: What is the fundamental physical principle that explains quantum correlations [^1] An equivalent way of phrasing this question is: given that there exist correlations satisfying relativity, but more Bell-non-local than the quantum ones, what in Nature prohibits their existence? Or even, why is Nature not more non-local? What limits Nature’s non-locality? All things considered, this program’s idea is to study the set of quantum correlations in Bell scenarios from the outside, investigating physically motivated device-independent restrictions strong enough to rule out, at best, all post-quantum behaviors.? Several principles to address this question, named as the problem of quantum correlations, have since then been proposed: Non-Triviality of Communication Complexity [12], Macroscopically Locality [13], Local Orthogonality [14] and Information Causality [15, 16] being examples of. All those trials focus on the problem of distant correlations, with the bipartite Bell scenario used as a playground for initial discussions, which could be further expanded to multipartite scenarios. In this work we have a special interest in the so-called Exclusivity Principle (EP) [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27], which avoids the necessity of parts, dealing with the problem of quantum correlations in Kochen-Specker (KS) contextuality scenarios [28].
@@
In 2010, Cabello, Severini, and Winter (CSW) presented an interesting connection between graph-theoretic concepts and KS-contextuality [29, 30]. This framework starts with a set of measurement events $e_{1},\ldots,e_{n}$. Two events are considered exclusive if they correspond to different outcomes of the same measurement [^2] A measurement event $e_{i}$ is to be understood as the performance of a given contextuality scenario, meaning $e_{i}=a,b,...,c|x,y,...,z$; in words, obtaining outcomes $a,b,...,c$ conditioned, respectively, on the chosen measurement context $x,y,...,z$. Two events $e=a,b,...,c|x,y,...,z$ and $e^{\prime}=a^{\prime},b^{\prime},...,c^{\prime}|x^{\prime},y^{\prime},...,z^{\prime}$ are said to be exclusive if $(x=x^{\prime}\land a\neq a^{\prime})\lor(y=y^{\prime}\land b\neq b^{\prime})\lor...\lor(z=z^{\prime}\land c\neq c^{\prime})$.. We associate each event $e_{i}$ with a vertex $i$, and we say that two vertices $i$ and $j$ are adjacent, denoted as $i\sim j$, if the corresponding measurement events $e_{i}$ and $e_{j}$ are exclusive. Based on these vertices and adjacency relations, we construct a graph $G$, which we refer to as the exclusivity graph.
@@
For a given exclusivity graph $G$, we consider theories that assign probabilities to the measurement events corresponding to its vertices. A behavior associated with $G$ is represented by a vector $(p(e_{1}),\ldots,p(e_{n}))$, where $p(e_{i})$ represents the probability of event $e_{i}$. Since any two exclusive events correspond to different outcomes of some measurement, the behavior of a graph must also satisfy the relation $p(e_{i})+p(e_{j})\leq 1$, whenever $i\sim j$ [31].
@@
The set of behaviors is determined by the exclusivity constraints defined by the graph $G$ and by the underlying physical theory used to describe the system. For instance, a behavior $(p(e_{1}),\ldots,p(e_{n}))$ is called deterministic noncontextual if it satisfies $p(e_{i})\in\{0,1\}$ and $p(e_{i})+p(e_{j})\leq 1$ for all pairs $i\sim j$ [32]. The set of all noncontextual (or classical) behaviors, denoted as $\mathrm{NC}(G)$, is then obtained by making probabilistic mixtures of such deterministic behaviors, a construction known as the convex hull of those deterministic noncontextual behaviors [^3] In graph-theoretical terms, this set is known as $\mathrm{STAB}(G)$ [30, 33].. Behaviors that do not belong to $\mathrm{NC}(G)$ are considered contextual.
@@
In quantum theory, the pairwise joint measurability of a set of observables implies the joint measurability of the entire set [34]. Consequently, for quantum behaviors, if a collection of measurement events ${e_{k}}$ is pairwise exclusive, the constraint $\sum_{k}p(e_{k})\leq 1$ must hold. The Exclusivity Principle (EP) extends this property as a universal hypothesis, proposing that this constraint should be satisfied in any physical theory. In other words, the EP asserts that for any set of pairwise exclusive measurement events ${e_{k}}$—that is, events that can be associated to mutually exclusive outcomes of a single mother measurement—the sum of their probabilities must satisfy $\sum_{k}p(e_{k})\leq 1$. We define EP1-behaviors as those that adhere to the Exclusivity Principle. The set of all such behaviors is referred to as the (single-copy) EP set [^5] In graph-theoretical terms, this set is referred to as $\mathrm{QSTAB}(G)$ [30, 33]., denoted by $E_{1}(G)$.
Cabello, Severini, and Winter [29, 30] showed that given any non-contextuality (NC) expression appropriately written as a positive linear combination of probabilities of measurement events, $S_{\omega}=\sum_{i}\omega_{i}p(e_{i})$, its corresponding classical (or non-contextual), quantum, and (single-copy) exclusivity bounds match three characteristic numbers of the associated weight-exclusivity graph, a weighted-graph created from the exclusivity graph associated to the events $e_{1},\ldots,e_{n}$, and the weight-vector $\omega=(\omega_{1},\ldots,\omega_{n})$. Specifically, this relationship is expressed as
$S_{\omega}\leq^{\mathrm{NC}}\alpha(G,\mathbf{\omega})\leq^{\mathrm{Q}}\vartheta(G,\mathbf{\omega})\leq^{\mathrm{E}_{1}}\alpha^{*}(G,\mathbf{\omega}),$ (1)
in other words, $\mathrm{max}_{\mathrm{NC}(G)}S_{\mathbf{\omega}}=\alpha(G,\mathbf{\omega})$, $\mathrm{max}_{\mathrm{Q}(G)}S_{\mathbf{\omega}}=\vartheta(G,\mathbf{\omega})$ and $\mathrm{max}_{\mathrm{E}_{1}(G)}S_{\mathbf{\omega}}=\alpha^{*}(G,\mathbf{\omega})$, where $\alpha(G,\mathbf{\omega})$ is the independence number, $\vartheta(G,\mathbf{\omega})$ is its Lovász number, and $\alpha^{*}(G,\mathbf{\omega})$ is its fractional packing number, being all these three numbers well-studied objects in graph theory (and combinatorial optimization) [35, 36, 37, 33, 31].
Now, since the quantum bound of any NC inequality corresponds to the Lovász number of its associated exclusivity graph $(G,\mathbf{\omega})$, and $\vartheta(G,\mathbf{\omega})$ is obtained from the set of quantum correlations of $G$, $\mathrm{Q(G)}$, if a physical principle is able to single out the latter (or, equivalently, justify why post-quantum graph behaviors are not allowed), then such principle corresponds to an answer of why is Nature not more contextual. This line of reasoning was first proposed in 2013 by Cabello [17], bringing the problem of quantum correlations to the CSW approach.
@@
By considering two copies of the pentagon graph (the exclusivity graph associated with the KCBS inequality, known for being the simplest NC inequality [38]), Cabello [17] was able to show that the maximum value allowed by EP to it is $\sqrt{5}$, which also equals its maximum quantum value. Shortly afterward, Yan proposed [21] a more general construction, involving the consideration of any given experiment and what is called its respective complementary experiment. This approach proved highly effective for applying the EP and enabled the extension of previous results. All results to be recalled in this paper, as well as the original ones to be shown, are derived from such underlying construction. Let us give it some special attention.
@@
Let $\{e_{i}\}$ be a set of $n$ measurement events with exclusivity graph $G$ and $\{e_{j}^{\prime}\}$ a set of also $n$ measurement events but with exclusivity graph $\bar{G}$, where $\bar{G}$ is the complementary graph [^6] Complementary graph $\bar{G}$ is defined with the same vertex set as $G$, i.e., $V(\bar{G})=V(G)$, but its edge set is defined by relation $ij\in E(\bar{G})\iff ij\notin E(G)$. of $G$. One then defines a set of composite events $\{f_{ij}=e_{i}\land e_{j}^{\prime}\}_{ij}$ ($f_{ij}$ occurs iff $e_{i}$ and $e_{j}^{\prime}$ occurs). The exclusivity graph encoding the exclusivity structure of events $f_{ij}$ is given by the disjunctive product [^7] Given graphs $G=(V(G),E(G))$ and $H=(V(H),E(H))$, $G*H$ is defined as: $V(G*H)=V(G)\times V(H)$; edge $(i,j^{\prime})(k,l^{\prime})\in E(G*H)$ iff $ik\in E(G)$ or $j^{\prime}l^{\prime}\in E(H)$. of $G$ and $\bar{G}$, namely $G*\bar{G}$. For any $G$, $\{f_{ii}\}_{i}$ corresponds to a subset of events in $G*\bar{G}$ consisting of only pairwise exclusive events. Consequently, EP implies
@@
At this point, it is further assumed that the experiments $G$ and $\bar{G}$ are completely independent, allowing the joint probabilities of local events to factorize as $p(e_{i},e_{i}^{\prime})=p(e_{i})p(e_{i}^{\prime})$. Surprisingly, even under this assumption of independence, one experiment can impose constraints on the other through the EP inequality, Ineq. (2). Such a paradigm is hereafter referred to as Yan’s construction. By making assumptions about the available behavior set of (or, equivalently, theory obeyed by) the complementary experiment and applying the resulting EP inequality, several significant results were obtained.
Yan first demonstrated [21] that, given $\mathrm{Q(\bar{G})}$, EP singles out the maximum quantum value of inequality $S=\sum_{i}p_{i}$ for experiment $G$. However, it is worth mentioning that there are infinitely many other inequalities [^8] Specifically, any inequality $S_{\omega}$ with weights differing from $1$ for each measurement event., which are based on the exclusivity structure of experiment $G$, for which Yan’s result says nothing about. In 2014, Amaral, Terra, and Cabello (ATC) established a stronger result [22]: given $\mathrm{Q(\bar{G})}$, EP singles out $\mathrm{Q(G)}$. This effectively resolves the limitation of the previous finding. Since EP picks out the entire quantum behavior set, it follows that for any chosen inequality $S_{\omega}$, EP also identifies its quantum maximum. Furthermore, ATC demonstrated that when the experiment consists of self-complementary graphs [^9] A graph $G$ is self-complementary if it is isomorphic to its complement $\bar{G}$; an example is the KCBS graph., EP forbids Post-Quantum sets [^10] Where by Post-Quantum sets we mean sets strictly larger than $\mathrm{Q(G)}$.. A logical next step would be to ask whether such a result holds even without the self-complementarity assumption, i.e., to show that EP forbids sets larger than the quantum for any given experiment. In 2019, Cabello demonstrated [26, 27] through a series of intricate arguments — for instance, considering an infinite number of copies of an experiment — that EP forbids sets larger than $\mathrm{Q(G)}$ for self-complementary graphs. A second key result in the same paper provides a systematic procedure for constructing a self-complementary graph $H(G)$ from any given initial graph $G$. By combining this graph construction technique with his first result and using some additional involved results from graph theory, Cabello was ultimately able to show that EP singles out $\mathrm{Q(G)}$ for any experiment. In some sense, the 2019 work solves the quantum correlations problem within the CSW approach. However, it does so by employing complex arguments, including constructions that extend far beyond those proposed by Yan.
In this work, we use Yan’s construction together with the notion of anti-blocking duality theory [39] to show that correlation sets strictly larger than $\mathrm{Q(G)}$ are ruled out by the EP.
@@ II Results
We begin by presenting the core technical contribution of this work in Proposition 1, where we establish that if the set of correlations for the complementary experiment is given by $X(\bar{G})$, then the EP constrains the maximum set of correlations for the original experiment to the anti-blocking set of $X(\bar{G})$. A direct consequence of this proposition, combined with a foundational graph-theoretical result [33], leads to our first result: assuming the set of correlations for the complementary experiment is the $\mathrm{E_{1}(G)}$, the EP restricts the set of correlations for the experiment to the classical set (Corollary 1). Lastly, we present the central result of this paper: quantum theory and EP exclude post-quantum theories. This can be seen as a generalization of ATC’s second result for any exclusivity graph. This generalization addresses the previously identified logical gap in the literature and provides valuable insights into the findings of the 2019 work.
@@
The above inequality can be seen as a half-space separation over the set of behaviors for the experiment $G$, where the probabilities $p(e_{i}^{\prime})$ are seen as the coefficients of this inequality. Moreover, we can progressively consider broader sets of achievable behaviors for the complementary experiment, and subsequently investigate the behavior sets of the experiment as defined by all resulting EP constraints of Ineq. (3). For the remainder of the paper, we will refer to these graph correlation sets as theories for experiments.
@@
Given that the theory describing the complementary experiment is $X(\bar{G})$, the largest theory allowed by EP for the experiment is $Y(G)=\mathrm{abl}X(\bar{G})$.
@@
On the one hand, $Y(G)$ is defined by all EP inequalities in form of Ineq. (3), which arise when considering all behaviors of the complementary experiment $\bar{G}$; this can be expressed mathematically as follows:
$Y(G)=\{p\geq 0\>|\>\sum_{i}p_{i}p_{i}^{\prime}\leq 1\>\forall\>p^{\prime}\in X(\bar{G})\}.$ (4)
@@
Straightforward comparison between Eq. (4) and Eq. (5) reveals that, under Yan’s construction, EP constraint Ineq. (3) matches exactly the defining property of an anti-blocking set. Therefore, $\mathrm{Y(G)}=\mathrm{abl}X(\bar{G})$. ∎
Although Proposition 1 holds for all theories $X(\bar{G})$, we will focus on the most relevant cases. These range from the smallest possible set to describe $\bar{G}$, corresponding to classical theory, $\mathrm{NC}(\bar{G})$, to the largest possible set consistent with EP, corresponding to a theory that only demands single-copy exclusivities, denoted $\mathrm{E_{1}}(\bar{G})$.
In the following, we investigate the restrictive power of EP in selecting theories for a given experiment, by assuming specific theories for the complementary experiment, including the limiting cases of classical and $E_{1}$ theories, as well as QT.
@@
By Proposition 1, the largest theory allowed by EP to the experiment is $Y(G)=\mathrm{abl\>}\mathrm{E_{1}(\bar{G})}$. However, as shown in Ref. [33], for any experiment, $E_{1}$ and classical theories are connected via the anti-blocking operation, specifically $\mathrm{E_{1}(\bar{G})}=\mathrm{abl\>}\mathrm{NC(G)}$. Consequently, it follows that $\mathrm{abl\>E_{1}(\bar{G})}=\mathrm{abl\>abl\>NC(G)}$. Since for any convex corner [^11] A set of vectors $V$ is a convex corner if it is nonempty, closed, convex, nonnegative and $0\leq\vec{u}\leq\vec{v}$ and $\vec{v}\in V\Rightarrow\vec{u}\in V$; see section $30$ of [33]. $D$, we know that $\mathrm{abl\>abl\>}D=D$ and both $\mathrm{E_{1}(\bar{G})}$ and $\mathrm{NC(G)}$ are convex corners [33], it results that $Y(G)=\mathrm{abl\>E_{1}(\bar{G})}=\mathrm{NC(G)}$. As for the second part, the distinction between the label of an experiment and its complement is entirely arbitrary. Therefore, by exchanging $G$ with $\bar{G}$, we obtain $\mathrm{E_{1}(G)}=\mathrm{abl\>NC(\bar{G})}$. ∎
The first part of Corollary 1 states that if EP holds and there exists a complementary experiment $\bar{G}$ where it is provable that all EP1-behaviors are accessible [^12] The case of interest being an imperfect graph; for perfect graphs $\mathrm{NC(\bar{G})=Q(\bar{G})=E_{1}(\bar{G})}$., then to the experiment $G$, only noncontextual behaviors are allowed. As for the second part, if EP is the only restriction and classical theory is accessible to $\bar{G}$, then to the experiment $G$, all EP1-behaviors are allowed. Notably, the larger the theory assumed for the complementary experiment, the stronger EP becomes in excluding behaviors from the experiment. Conversely, the smaller the theory assumed for the complementary experiment, the weaker EP becomes in excluding behaviors from the experiment. This mechanism is precisely what underpins the proof of our main result. Before showing it, we remark that Proposition 1 recovers result $1$ due to ATC in Ref. [22].
@@
Given $\mathrm{Q(\bar{G})}$, EP singles out $\mathrm{Q(G)}$.

### v2

@@ I Introduction
Quantum Theory (QT) is one of the most successful scientific theories ever created by humankind; in that its predictions match with great accuracy empirical data. One paramount example is the loophole-free confirmation of the existence of quantum Bell-nonlocality [Giustina2015, Hensen2015, Shalm2015, Hensen2016, Li2018]. Nonetheless, it remains a challenge to formulate it as a consequence of physically justifiable principles (of Nature), similar to what occurs with Einstein’s Special Relativity. Several distinct approaches have been proposed to address this problem. For example, certain approaches adopt an axiomatic framework, beginning with a more physically intuitive set of axioms from which one aims to derive the standard and abstract postulates of quantum theory (the Hilbert space structure, density matrices and measurement operators acting on it, and Born’s rule) [hardy2001quantumtheoryreasonableaxioms, hardy2011reformulatingreconstructingquantumtheory, Masanes2011, Chiribella2010, Chiribella2011]. Another program is the one started by Popescu and Rohrlich’s seminal work in trying to answer the question: are quantum correlations singled out by the non-signaling principle [Popescu1994]? The answer was found to be negative, as there exist correlations that satisfy the non-signaling condition but cannot be reproduced within the framework of quantum theory, with Popescu-Rohrlich boxes being the most prominent example [Brunner2014]. Although this approach does not fully explain quantum correlations, it raises the broader question: What is the fundamental physical principle that explains quantum correlations [^1] An equivalent way of phrasing this question is: given that there exist correlations satisfying relativity, but more Bell-non-local than the quantum ones, what in Nature prohibits their existence? Or even, why is Nature not more non-local? What limits Nature’s non-locality? All things considered, this program’s idea is to study the set of quantum correlations in Bell scenarios from the outside, investigating physically motivated device-independent restrictions strong enough to rule out, at best, all post-quantum behaviors.? Several principles to address this question, named as the problem of quantum correlations, have since then been proposed: Non-Triviality of Communication Complexity [Brassard2006], Macroscopically Locality [Navascus2009], Local Orthogonality [Fritz2013] and Information Causality [Pawowski2009, Pollyceno2023] being examples of such principles. All those trials focus on the problem of distant correlations, with the bipartite Bell scenario used as a playground for initial discussions, which could be further expanded to multipartite scenarios. In this work we have a special interest in the so-called Exclusivity Principle (EP) [Cabello2013a, cabello2013proposedexperimentexcludehigherthanquantum, nawareg2013bounding, Cabello2013b, Yan2013, Amaral2014, Cabello2014, Cabello2015, jia2022exclusivity, Cabello2019, Cabello2019b], which avoids the necessity of parts, dealing with the problem of quantum correlations in Kochen-Specker (KS) contextuality scenarios [Budroni2022].
@@
In 2010, Cabello, Severini, and Winter (CSW) presented an interesting connection between graph-theoretic concepts and KS-contextuality [https://doi.org/10.48550/arxiv.1010.2163, Cabello2014c]. This framework starts with a set of measurement events $e_{1},\ldots,e_{n}$. Two events are considered exclusive if they correspond to different outcomes of the same measurement [^2] A measurement event $e_{i}$ is to be understood as the performance of a given contextuality scenario, meaning $e_{i}=a,b,...,c|x,y,...,z$; in words, obtaining outcomes a,b,,ca,b,...,c conditioned, respectively, on the chosen measurement context x,y,,zx,y,...,z. Two events e=a,b,,c|x,y,,ze=a,b,...,c|x,y,...,z and $e^{\prime}=a^{\prime},b^{\prime},...,c^{\prime}|x^{\prime},y^{\prime},...,z^{\prime}$ are said to be exclusive if $(x=x^{\prime}\land a\neq a^{\prime})\lor(y=y^{\prime}\land b\neq b^{\prime})\lor...\lor(z=z^{\prime}\land c\neq c^{\prime})$.. We associate each event $e_{i}$ with a vertex $i$, and we say that two vertices $i$ and $j$ are adjacent, denoted as $i\sim j$, if the corresponding measurement events $e_{i}$ and $e_{j}$ are exclusive. Based on these vertices and adjacency relations, we construct a graph $G$, which we refer to as the exclusivity graph.
@@
For a given exclusivity graph $G$, we consider theories that assign probabilities to the measurement events corresponding to its vertices. A behavior associated with $G$ is represented by a vector $(p(e_{1}),\ldots,p(e_{n}))$, where $p(e_{i})$ represents the probability of event $e_{i}$. Since any two exclusive events correspond to different outcomes of some measurement, the behavior of a graph must also satisfy the relation $p(e_{i})+p(e_{j})\leq 1$, whenever $i\sim j$ [amaral_graph_2018].
@@
The set of behaviors is determined by the exclusivity constraints defined by the graph $G$ and by the underlying physical theory used to describe the system. For instance, a behavior $(p(e_{1}),\ldots,p(e_{n}))$ is called deterministic noncontextual if it satisfies $p(e_{i})\in\{0,1\}$ and $p(e_{i})+p(e_{j})\leq 1$ for all pairs $i\sim j$ [bharti_robust_2019]. The set of all noncontextual (or classical) behaviors, denoted as $\mathrm{NC(G)}$, is then obtained by making probabilistic mixtures of such deterministic behaviors, a construction known as the convex hull of those deterministic noncontextual behaviors [^3] In graph-theoretical terms, this set is known as $\mathrm{STAB}(G)$ [Cabello2014c, Knuth1994].. Behaviors that do not belong to $\mathrm{NC(G)}$ are considered contextual.
@@
[…]
@@
By considering two copies of the pentagon graph (the exclusivity graph associated with the KCBS inequality, known for being the simplest NC inequality [Klyachko2008]), Cabello [Cabello2013a] was able to show that the maximum value allowed by EP to it is $\sqrt{5}$, which also equals its maximum quantum value. Shortly afterward, Yan proposed [Yan2013] a more general construction, involving the consideration of any given experiment and what is called its respective complementary experiment. This approach proved highly effective for applying the EP and enabled the extension of previous results. All results to be recalled in this paper, as well as the original ones to be shown, are derived from such underlying construction. Let us give it some special attention.
@@
Let $\{e_{i}\}$ be a set of $n$ measurement events with exclusivity graph $G$ and $\{e_{j}^{\prime}\}$ a set of also $n$ measurement events but with exclusivity graph $\bar{G}$, where $\bar{G}$ is the complementary graph [^7] Complementary graph $\bar{G}$ is defined with the same vertex set as $G$, i.e., $V(\bar{G})=V(G)$, but its edge set is defined by relation $ij\in E(\bar{G})\iff ij\notin E(G)$. of $G$. One then defines a set of composite events $\{f_{ij}=e_{i}\land e_{j}^{\prime}\}_{ij}$ ($f_{ij}$ occurs iff $e_{i}$ and $e_{j}^{\prime}$ occurs). The exclusivity graph encoding the exclusivity structure of events $f_{ij}$ is given by the disjunctive product [^8] Given graphs $G=(V(G),E(G))$ and $H=(V(H),E(H))$, GHG*H is defined as: $V(G*H)=V(G)\times V(H)$; edge $(i,j^{\prime})(k,l^{\prime})\in E(G*H)$ iff $ik\in E(G)$ or $j^{\prime}l^{\prime}\in E(H)$. of $G$ and $\bar{G}$, namely $G*\bar{G}$. For any $G$, $\{f_{ii}\}_{i}$ corresponds to a subset of events in $G*\bar{G}$ consisting of only pairwise exclusive events. Consequently, the EP implies
@@
At this point, it is further assumed that the experiments $G$ and $\bar{G}$ are completely independent, allowing the joint probabilities of local events to factorize as $p(e_{i},e_{i}^{\prime})=p(e_{i})p(e_{i}^{\prime})$. Surprisingly, even under this assumption of independence, one experiment can impose constraints on the other through the EP inequality, Ineq. (2). Such a paradigm is hereafter referred to as Yan’s construction. By making assumptions about the available behavior set of the complementary experiment and applying the resulting EP inequality, several significant results were obtained.
Yan first demonstrated [Yan2013] that, given $\mathrm{Q(\bar{G})}$, EP singles out the maximum quantum value of inequality $S=\sum_{i}p_{i}$ for experiment $G$. However, it is worth mentioning that there are infinitely many other inequalities [^9] Specifically, any inequality $S$ with weights differing from $1$ for each measurement event., which are based on the exclusivity structure of experiment $G$, for which Yan’s result says nothing about. In 2014, Amaral, Terra, and Cabello (ATC) established a stronger result [Amaral2014]: given $\mathrm{Q(\bar{G})}$, EP singles out $\mathrm{Q(G)}$. This effectively resolves the limitation of the previous finding. Since EP picks out the entire quantum behavior set, it follows that for any chosen inequality $S$, EP also identifies its quantum maximum. Furthermore, ATC demonstrated that when the experiment consists of self-complementary graphs [^10] A graph $G$ is self-complementary if it is isomorphic to its complement $\bar{G}$; an example is the KCBS graph., EP forbids Post-Quantum sets [^11] Where by Post-Quantum sets we mean sets strictly larger than $\mathrm{Q(G)}$.. A logical next step would be to ask whether such a result holds even without the self-complementarity assumption, i.e., to show that EP forbids sets larger than the quantum for any given experiment. In 2019, Cabello demonstrated [Cabello2019, Cabello2019b] through a series of intricate arguments — for instance, considering an infinite number of copies of an experiment — that EP forbids sets larger than $\mathrm{Q(G)}$ for self-complementary graphs. A second key result in the same paper provides a systematic procedure for constructing a self-complementary graph $H(G)$ from any given initial graph $G$. By combining this graph construction technique with his first result and using some additional involved results from graph theory, Cabello was ultimately able to show that EP singles out $\mathrm{Q(G)}$ for any experiment. In some sense, the 2019 work solves the quantum correlations problem within the CSW approach. However, it does so by employing complex arguments, including constructions that extend far beyond those proposed by Yan.
In this work, we use Yan’s construction together with the notion of anti-blocking theory [Fulkerson1972] to show that correlation sets strictly larger than $\mathrm{Q(G)}$ are ruled out by the EP.
@@ II Results
We begin by presenting the core technical contribution of this work in Proposition 1, where we establish that if the set of correlations for the complementary experiment is given by $\mathrm{X(\bar{G})}$, then the EP constrains the maximum set of correlations for the original experiment to the anti-blocking set of $\mathrm{X(\bar{G})}$. A direct consequence of this proposition, combined with a foundational graph-theoretical result [Knuth1994], leads to our first result: assuming the set of correlations for the complementary experiment is the $\mathrm{E_{1}(\bar{G})}$, the EP restricts the set of correlations for the experiment to the classical set (Corollary 1). Lastly, we present the central result of this paper: assuming that Nature is capable of realizing every behavior allowed by quantum theory and that the EP holds, it follows that post-quantum correlations are forbidden. This can be seen as a generalization of ATC’s second result to any exclusivity graph. In doing so, our work addresses a methodological convolutedness in the literature, where the 2019 study deviated from the established trend of utilizing Yan’s construction in favor of a significantly more complex approach. By employing Yan’s framework directly, we pave this road and demonstrate that results similar to those in the 2019 work can be obtained through a much simpler methodology, offering additional insights into its findings.
@@
The above inequality can be seen as a half-space separation over the set of behaviors for the experiment $G$, where the probabilities $p(e_{i}^{\prime})$ are seen as the coefficients of this inequality. Moreover, we can progressively consider broader sets of achievable behaviors for the complementary experiment, and subsequently investigate the behavior sets of the experiment as defined by all resulting EP constraints of Ineq. (3).
@@
Given that the correlation set describing the complementary experiment is $\mathrm{X(\bar{G})}$, the largest correlation set allowed by the EP for the experiment is $\mathrm{Y(G)}=\mathrm{abl}\>\mathrm{X(\bar{G})}$.
@@
On the one hand, $\mathrm{Y(G)}$ is defined by all EP inequalities in form of Ineq. (3), which arise when considering all behaviors of the complementary experiment $\bar{G}$; this can be expressed mathematically as follows:
$\mathrm{Y(G)}=\{p\geq 0\>|\>\sum_{i}p_{i}p_{i}^{\prime}\leq 1\>\forall\>p^{\prime}\in\mathrm{X(\bar{G})}\}.$ (4)
@@
Straightforward comparison between Eq. (4) and Eq. (5) reveals that, under Yan’s construction, EP constraint Ineq. (3) matches exactly the defining property of an anti-blocking set. Therefore, $\mathrm{Y(G)}=\mathrm{abl}\>\mathrm{X(\bar{G})}$. ∎
Although Proposition 1 holds for all sets $\mathrm{X(\bar{G})}$, we will focus on the most relevant cases. These range from the smallest possible set to describe $\bar{G}$, corresponding to classical (or noncontextual) behaviors, $\mathrm{NC(\bar{G})}$, to the largest possible set consistent with the EP, corresponding to correlations that are only required to satisfy single-copy exclusivities, denoted $\mathrm{E_{1}}(\bar{G})$.
In the following, we investigate the restrictive power of EP in selecting correlation sets for a given experiment, by assuming specific sets for the complementary experiment, including the limiting cases of classical and $E_{1}$ correlations, as well as the quantum one.
@@
By Proposition 1, the largest correlation set allowed by the EP to the experiment is $\mathrm{Y(G)}=\mathrm{abl\>}\mathrm{E_{1}(\bar{G})}$. However, as shown in Ref. [Knuth1994], for any experiment, $E_{1}$ and classical correlation sets are connected via the anti-blocking operation, specifically $\mathrm{E_{1}(\bar{G})}=\mathrm{abl\>}\mathrm{NC(G)}$. Consequently, it follows that $\mathrm{abl\>E_{1}(\bar{G})}=\mathrm{abl\>abl\>NC(G)}$. Since for any convex corner [^12] A set of vectors $V$ is a convex corner if it is nonempty, closed, convex, nonnegative and $0\leq\vec{u}\leq\vec{v}$ and $\vec{v}\in V\Rightarrow\vec{u}\in V$; see section $30$ of [Knuth1994]. $D$, we know that $\mathrm{abl\>abl\>}D=D$ and both $\mathrm{E_{1}(\bar{G})}$ and $\mathrm{NC(G)}$ are convex corners [Knuth1994], it results that $\mathrm{Y(G)}=\mathrm{abl\>E_{1}(\bar{G})}=\mathrm{NC(G)}$. As for the second part, the distinction between the label of an experiment and its complement is entirely arbitrary. Therefore, by exchanging $G$ with $\bar{G}$, we obtain $\mathrm{E_{1}(G)}=\mathrm{abl\>NC(\bar{G})}$. ∎
The first part of Corollary 1 states that if EP holds and there exists a complementary experiment $\bar{G}$ where it is provable that all EP1-behaviors are accessible [^13] The case of interest being an imperfect graph; for perfect graphs $\mathrm{NC(\bar{G})=Q(\bar{G})=E_{1}(\bar{G})}$., then to the experiment $G$, only noncontextual behaviors are allowed. As for the second part, if the EP is the only restriction and only classical behaviors are accessible to $\bar{G}$, then to the experiment $G$, all EP1-behaviors are allowed. Notably, the larger the correlation set assumed for the complementary experiment, the stronger EP becomes in excluding behaviors from the experiment. Conversely, the smaller the correlation set assumed for the complementary experiment, the weaker EP becomes in excluding behaviors from the experiment. This mechanism is precisely what underpins the proof of our main result. Before showing it, we remark that Proposition 1 recovers result $1$ due to ATC in Ref. [Amaral2014].
@@
(Result 1 of Ref. [Amaral2014]) Given $\mathrm{Q(\bar{G})}$, EP singles out $\mathrm{Q(G)}$.